OpenStudy (anonymous):

MEDAL!:) Find the coordinates of the center of the given circle. (x - 5)2 + (y + 3)2 = 25 (5, 3) (5, -3) (-5, 3)

OpenStudy (anonymous):

\[\left( x-h \right)^{2}+\left( y-k \right)^{2}=r ^{2}\] centre is (h,k) and radius r.